OpenStudy (anonymous):

Victor calculates the percentage interest rate he needs to make $20 in simple annual interest over 2 months on a deposit of $1000 . Where did he go wrong?

OpenStudy (campbell_st):

\[20 = 1000\times r \times(1/6) \] \[r = \frac{20 \times 6}{1000}\] multiply r by 100 to get the percentage